OpenStudy (fools101):

How would i do Find the zeros of the following polynomial functions. For the 4th degree polynomials the first step of factoring is provided. Factor completely and solve y = 2x4 + 16x3 + 32x2 = 2x2(x2 + 8x + 16)

OpenStudy (fools101):

help please!

jimthompson5910 (jim_thompson5910):

2x^2(x^2 + 8x + 16) = 0 2x^2 = 0 or x^2 + 8x + 16 = 0 2x^2 = 0 or (x+4)^2 = 0 x^2 = 0 or (x+4)^2 = 0 x = 0 or x+4 = 0 x = 0 or x = -4

OpenStudy (fools101):

THANK U!!!